OpenStudy (anonymous):

One ship is sailing south at a rate of 5 knots, and another is sailing east at a rate of 10 knots. At 2 P.M. the second ship was at the place occupied by the first ship one hour before. at what time was the distance between the ships not changing.?

OpenStudy (anonymous):

Did you draw a picture yet. i find this really helps.

OpenStudy (anonymous):

@ChmE 

OpenStudy (anonymous):

@ChmE ??

OpenStudy (anonymous):

So let \(t\) be our variable, and it will represent hours.

OpenStudy (anonymous):

what happen next? @wio

OpenStudy (anonymous):

Let f(t) be the position of ship sailing south, and g(t) be the position ship sailing east.

OpenStudy (anonymous):

f(0) = g(2)

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

I'm interested in the sol'n as well. How did you come up with f(0) = g(2)

OpenStudy (anonymous):

I thought it would be one hour

OpenStudy (anonymous):

You're right, it's f(0) = g(1)

OpenStudy (anonymous):

@bii17 sry for not responding earlier. I wasn't confident in my sol'n and didn't want to possibly steer you down a wrong path at the risk of confusing you.

OpenStudy (anonymous):

ChmE, what is your solution?

OpenStudy (anonymous):

@ChmE that's ok.. @wio can i see your solution.. i really dont get the problem..

OpenStudy (anonymous):

My solution is to find the position functions, then use the Pythagorean theorem on them to get a 'change in distance' function. Then take the derivative. Then find when that is 0

OpenStudy (anonymous):

can you show it step by step ?? @wio

OpenStudy (anonymous):

Okay bii17. What is the anti-derivative of 5?

OpenStudy (anonymous):

anti derivative??? @wio

OpenStudy (anonymous):

I don't understand what the question is asking. Is it looking for at what time is the distance the same as the initial change in distance? Thats how I read it

OpenStudy (anonymous):

Yes, what does \(\int5dx\) equal?

OpenStudy (anonymous):

we havent discussed it in our class.. no idea

OpenStudy (anonymous):

Hmmmm, okay, well let's say that the ship started as position 0, it is moving 5 knots, how many knot-hours has it traveled in t hours?

OpenStudy (anonymous):

Let a knot-hour be the distance you travel in an hour at the speed of a knot. Our ship will have traveled 5t

OpenStudy (anonymous):

Since distance = speed * time.

OpenStudy (anonymous):

So f(t) = 5t. Now g(t) = 2t + c. Where c is the initial position. We don't know it yet, but we can use f(0) = g(1) to find out c.

OpenStudy (anonymous):

5(0) = 2(1) + c => c = -2

OpenStudy (anonymous):

So g(t) = 2t-2 Are you following?

OpenStudy (anonymous):

@wio yes..

OpenStudy (anonymous):

Now, since they are traveling perpendicular to each other... how can we find the distance between them with respect to time?

OpenStudy (anonymous):